OpenStudy (anonymous):

solve the equation x(square) -4x +1 =0

OpenStudy (anonymous):

use quadratic formula, you'll get: \[x=2+/-2\sqrt{3}\]

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

\[x^2 - 4x + 1 = 0\] \[x= 2 +/ - 2\sqrt{3}\]

OpenStudy (anonymous):

need detailed solution

OpenStudy (bobbyleary):

\[ax ^{2} + bx + c = 0\] To solve, you need to use the quadratic equation: \[x = (-b \pm \sqrt{b ^{2}-4ac}) / 2a\]"2a" divides both the "-b" and the square root. In your problem you have a = 1, b = -4, and c = 1 \[x = (-(-4)\pm \sqrt{(-4)^{2}-4(1)(1)})/2(1)\] Simplify to get...\[x = (4/2) \pm (\sqrt{12}/2) = 2 \pm (2\sqrt{3}/2) = 2 \pm \sqrt{3}\]